OpenStudy (anonymous):

Write the expression as a complex number in standard form: (5-7i)(8+6i) Please show the steps :)

OpenStudy (saifoo.khan):

82 - 26 i sorry my calc dont show the steps.

OpenStudy (anonymous):

YAY! That's what I got, I wanted to make sure I was right. Thank you!

OpenStudy (anonymous):

(5-7i)(8+6i) =8(5-7i)+6i(5-7i) can you do the next step?

OpenStudy (anonymous):

Ummmmmmmm, will I get the same answer as above?

OpenStudy (anonymous):

note i^2=-1

OpenStudy (anonymous):

yes the same answer

OpenStudy (anonymous):

Ok, I think I'll stick with the method that I have, thank you so much though! :)

OpenStudy (anonymous):

ok, if you have a calculator that have complex no. it will be easier..lol

OpenStudy (anonymous):

=40-56i+30i-42i^2 now collecting like or same terms

OpenStudy (anonymous):

82 - 26 i answer

OpenStudy (anonymous):

good luck now, thnx